# The Negation Of Singer’s Conjecture For The Sixth Algebraic Transfer

### v1

@@ 1 Introduction and statement of the main outcome
    $\mathcal{S}_{q}(n)=\sum_{\begin{subarray}{c}(c_{0},c_{1},\dots)\in\mathbb{Z}_{\geq 0},\\ \sum c_{m}=q,\,\sum c_{m}2^{m}=q+n\end{subarray}}\frac{q!}{\prod_{m\geq 0}c_{m}!}.$
@@
Let $[(QP_{q})_{n}]^{GL(q)}$ denote the subspace of $(QP_{q})_{n}$ comprising all $GL(q)$-invariants of degree $n.$ Consequently, the domain of the algebraic transfer is dual to the invariant $[(QP_{q})_{n}]^{GL(q)}$ for any $n.$ It should be noted that the bi-graded sum $\bigoplus_{q,n}[(QP_{q})_{n}]^{GL(q)}$ possesses a co-algebra structure. (This fact is derived from the co-algebra structure on $\bigoplus_{q}H^{*}(\mathcal{V}^{q}),$ which comes from the natural isomorphisms $H^{*}(\mathcal{V}^{q})\cong H^{*}(\mathcal{V}^{i})\otimes_{\mathbb{F}_{2}}H^{*}(\mathcal{V}^{j})$ with $i+j=q.$) Therefore, dualizing the co-algebra yields an algebraic structure on the domain of the total transfer $\mathsf{Tr}_{*}$, as previously mentioned.
@@
The conjecture is also very difficult to attack, partly because the calculation of both the domain and the codomain of $Tr_{q}(\mathbb{F}_{2})$ is not easy. It has remained an open problem for over three decades when $q\geq 4.$ Our recent work, as presented in [22, 23, 24], has successfully confirmed the conjecture’s validity for $q=4.$ In this paper, we show that the conjecture does not remain valid for the $q=6$ case.
@@
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [34], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand and, unfortunately, its correctness has not yet been verified using a computer algebra system (e.g., OSCAR, Magma, or SageMath). In fact, verifying Sum’s counterexample [34] on a computer algebra system is also very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [28]). The difficulty in verification is precisely why Conjecture 1.1 had effectively remained open for all homology degrees $q\geq 5$ until we now refute it by Theorems 1.2, 1.4 and Corollary 1.3 for the case $(q,n)=(6,36),$ and the result have been fully verified using computer algebra systems.
@@ 2 A few preliminaries
$g\equiv\beta(x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}+h_{1}+h_{2}+h_{3}+h_{4}+h_{5}).$
@@
In the next step, we will explicitly compute the basis for ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ and the $GL(6)$-invariant $[{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)}.$ As mentioned above, computing these spaces by hand seems infeasible and error-prone due to the prohibitively large number of input monomials (specifically, by the formula in [28], $\dim(P_{6})_{36}=\binom{36+(6-1)}{6-1}=749398$). To overcome this difficulty, we will construct an algorithmic program implemented in the computer algebra system OSCAR [39] that allows us to explicitly determine the basis of the spaces ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,n)})$ and of their $GL(q)$-invariant subspaces for any $q$ and $n$ satisfying $n-q$ even. Based on the previously obtained results for the $(GL(q)$-invariants of the target space of the Kameko homomorphism $(\widetilde{Sq}^{0}_{*})_{(q,n)}$ (computed via our algorithm in [29]), we proceed as follows: We first compute the $(GL(q)$-invariants in degree $(n-q)/2$ for the target of $(\widetilde{Sq}^{0}_{*})_{(q,n)}$ by the method of [29]. These target invariants are then used as seeds: we apply the inverse Kameko lift $\psi:\ (P_{q})_{\frac{n-q}{2}}\longrightarrow(P_{q})_{n},\ \ x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}\longmapsto x_{1}^{2e_{1}+1}\cdots x_{q}^{2e_{q}+1},$ and, on the subset of admissible coordinates contained in ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,n)})$, we solve the stacked linear systems enforcing $(\rho_{j}-\mathrm{Id})\,f=0\quad(j=1,\dots,q-1)$ and $(\rho_{q}-\mathrm{Id})\,f=0,$ thereby correcting the lifts to genuine $GL(q)$-invariants. In parallel, within the kernel itself we perform a weightwise computation of $\Sigma_{q}$- and $GL(q)$-invariants and then apply a largest-weight correction to non-zero $GL(q)$-invariants. Below we construct in detail our algorithm as sketched.
$\displaystyle\trianglerightIntegers$ $q\geq 1$, $d\geq 0$ with $d\equiv q\pmod{2}$; base field $\mathbb{F}_{2}$.
@@
$\displaystyle\trianglerightfunction$ $\Sigma/GL$-On-Kernel-Weight($\mathsf{DS},\;\ker L,\;\mathcal{I}_{w}$)
@@
$\displaystyle\triangleright$  (S-stage) Build stacked matrix of $(\rho_{j}-\mathrm{Id})\sum_{i}\gamma_{i}u_{i}$ for $j=1,\dots,q-1$; find nullspace
$\displaystyle\triangleright$  Obtain $\Sigma$-basis $\{\sum_{i}\gamma^{(t)}_{i}u_{i}\}_{t}$
$\displaystyle\triangleright$  (GL-stage) Build matrix of $(\rho_{q}-\mathrm{Id})\sum_{t}\beta_{t}(\Sigma\text{-basis})_{t}$; find nullspace
@@
$\displaystyle\triangleright$  return (S-basis, GL-basis) in this weight
@@
$\displaystyle\triangleright$   $(\text{Sigma}[\omega],\text{GL}[\omega])\leftarrow$ $\Sigma/GL$-On-Kernel-Weight($\mathsf{DS\_src},\ker L,\mathcal{I}_{\omega}$)
@@
$\displaystyle\triangleright$  [Step 4] Correction inside kernel (largest weight with $GL\neq 0$):
$\displaystyle\triangleright$  if all $\text{GL}[\omega]$ are empty then report $GL(q)$-invariants in kernel $=0$
@@
$\displaystyle\triangleright$   pick $\omega^{\star}=\max\{\omega:\text{GL}[\omega]\neq\emptyset\}$; set $\mathcal{L}:=\{i\in\mathcal{K}:\text{weight}(i)<\omega^{\star}\}$
$\displaystyle\triangleright$   for each $g_{\max}\in\text{GL}[\omega^{\star}]$ do
@@
    Bit-packed matrices and Gaussian elimination over $\mathbb{F}_{2}$. All nullspace computations (Kameko matrix, $\Sigma$-stage, $GL$-stage) use bit-packed matrices, so elimination and back-substitution become word-level XOR operations. This is cache-friendly and significantly faster than dense arithmetic over $\mathbb{F}_{2}$.

### v2

@@ 1 Introduction and statement of the main outcome
    (*) $\mathcal{S}_{q}(n)=\sum_{\begin{subarray}{c}(c_{0},c_{1},\dots)\in\mathbb{Z}_{\geq 0},\\ \sum c_{m}=q,\,\sum c_{m}2^{m}=q+n\end{subarray}}\frac{q!}{\prod_{m\geq 0}c_{m}!}.$
@@
To provide readers with the context related to the inequality above, we restate the following important fact, which has been mentioned in [31, Remark 2.3]: A key element in analyzing $(QP_{q})_{n}$ is the set of ”spike” monomials—monomials where every exponent is of the form $2^{m}-1$. A classical result establishes that spike monomials do not belong to $\mathscr{A}^{>0}\cdot P_{q}$ (see also [37]). Additionally, in [18], Mothebe constructed a rather involved recursive function to enumerate all spike monomials of degree $n=2^{q-1}-q$, namely
(**) $B(q,2^{q-1}-q)$ $=q\cdot B\!\left(q-1,\,2^{q-2}-(q-1)\right)+\binom{q}{3}\,B\!\left(q-3,\,2^{q-4}-(q-3)\right)$
$\quad+\sum_{5\leq r\leq q-2}\biggl{[}\sum_{(b_{1},b_{2},\ldots)\in[S^{r}(q)]}\binom{q}{q-r,\,b_{1},\,b_{2},\,\ldots}\biggr{]}B\!\left(q-r,\,2^{q-(r+1)}-(q-r)\right)$
$\quad+\sum_{(b_{1},b_{2},\ldots)\in[S^{q-1}(q)]}\binom{q}{\,b_{1}+1,\,b_{2},\,b_{3},\,\ldots}\!$,
where $[S^{r}(q)]$ is a family generated by certain tree constructions (see [18] for details), and $B(q,2^{q-1}-q)$ denotes the number of spikes of degree $2^{q-1}-q$. However, (**1) addresses only the special degree $2^{q-1}-q$. In the formula (*1), we derive an explicit general formula that applies to arbitrary $q$ and $n$.
For comparison, our formula (*1) reproduces the values given in [18] via Mothebe’s recursion (**1):
$\mathcal{S}_{2}(0)=B(2,0)=1,\qquad\mathcal{S}_{3}(1)=B(3,1)=3,\qquad\mathcal{S}_{4}(4)=B(4,4)=13,$
$\mathcal{S}_{5}(11)=B(5,11)=75,\qquad\mathcal{S}_{6}(26)=B(6,26)=525,\qquad\mathcal{S}_{7}(57)=B(7,57)=4,347.$
Remarkably, Mothebe [18] gives a worked example for $q=11$ and $n=1013$ with the claim $B(11,1013)=135,029,697$. Nevertheless, this hand computation is inaccurate. In [27], we implemented a SageMath algorithm based on Mothebe’s method and obtained $B(11,1013)=68,958,747$. Applying our closed form (*1) independently yields the same value, $\mathcal{S}_{11}(1013)=68,958,747$. This shows that the hand calculation in [18] for $q=11$, $n=1013$ is not true.
Let now $[(QP_{q})_{n}]^{GL(q)}$ denote the subspace of $(QP_{q})_{n}$ comprising all $GL(q)$-invariants of degree $n.$ Consequently, the domain of the algebraic transfer is dual to the invariant $[(QP_{q})_{n}]^{GL(q)}$ for any $n.$ It should be noted that the bi-graded sum $\bigoplus_{q,n}[(QP_{q})_{n}]^{GL(q)}$ possesses a co-algebra structure. (This fact is derived from the co-algebra structure on $\bigoplus_{q}H^{*}(\mathcal{V}^{q}),$ which comes from the natural isomorphisms $H^{*}(\mathcal{V}^{q})\cong H^{*}(\mathcal{V}^{i})\otimes_{\mathbb{F}_{2}}H^{*}(\mathcal{V}^{j})$ with $i+j=q.$) Therefore, dualizing the co-algebra yields an algebraic structure on the domain of the total transfer $\mathsf{Tr}_{*}$, as previously mentioned.
@@
The conjecture is also very difficult to attack, partly because the calculation of both the domain and the codomain of $Tr_{q}(\mathbb{F}_{2})$ is not easy. It has remained an open problem for over three decades when $q\geq 4.$ Our recent works, as presented in [23, 24, 25], have successfully confirmed the conjecture’s validity for $q=4.$ In this paper, we show that the conjecture does not remain valid for the $q=6$ case.
@@
Recently, a result for the case $q=5$ has been proposed in a preprint by Nguyen Sum [35], which provides a counterexample to Conjecture 1.1 in bidegree $(5,5+108).$ This result was computed entirely by hand and, unfortunately, its correctness has not yet been verified using a computer algebra system (e.g., OSCAR, Magma, or SageMath). In fact, verifying Sum’s counterexample [35] on a computer algebra system is also very difficult, as the degree $n=108$ results in a prohibitively large number of input monomials for $(q,n)=(5,108),$ namely $\binom{108+(5-1)}{5-1}=6,210,820.$ (The general formula for this calculation, $\binom{n+q-1}{q-1}$, is given in our recent work [29]). The difficulty in verification is precisely why Conjecture 1.1 had effectively remained open for all homology degrees $q\geq 5$ until we now refute it by Theorems 1.2, 1.4 and Corollary 1.3 for the case $(q,n)=(6,36),$ and the result has been fully verified using computer algebra systems.
@@ 2 A few preliminaries
We refer the reader to our work [30] for a detailed proof of this result.
@@
$g\equiv\beta\bigg{(}x_{1}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{4}x_{6}^{4}+\sum_{1\leq j\leq 5}h_{j}\bigg{)},\,\beta\in\mathbb{F}_{2}.$
@@
In the next step, we will explicitly compute the basis for ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ and the $GL(6)$-invariant $[{\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})]^{GL(6)}.$ As mentioned above, computing these spaces by hand seems infeasible and error-prone due to the prohibitively large number of input monomials (specifically, by the formula in [29], $\dim(P_{6})_{36}=\binom{36+(6-1)}{6-1}=749398$). To overcome this difficulty, we will construct an algorithmic program implemented in the computer algebra system OSCAR [40] that allows us to explicitly determine the basis of the spaces ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,n)})$ and of their $GL(q)$-invariant subspaces for any $q$ and $n$ satisfying $n-q$ even. Based on the previously obtained results for the $(GL(q)$-invariants of the target space of the Kameko homomorphism $(\widetilde{Sq}^{0}_{*})_{(q,n)}$ (computed via our algorithm in [30]), we proceed as follows: We first compute the $(GL(q)$-invariants in degree $(n-q)/2$ for the target of $(\widetilde{Sq}^{0}_{*})_{(q,n)}$ by the method of [30]. These target invariants are then used as seeds: we apply the inverse Kameko lift $\psi:\ (P_{q})_{\frac{n-q}{2}}\longrightarrow(P_{q})_{n},\ \ x_{1}^{e_{1}}\cdots x_{q}^{e_{q}}\longmapsto x_{1}^{2e_{1}+1}\cdots x_{q}^{2e_{q}+1},$ and, on the subset of admissible coordinates contained in ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(q,n)})$, we solve the stacked linear systems enforcing $(\rho_{j}-\mathrm{Id})\,f\equiv 0\quad(j=1,\dots,q-1)$ and $(\rho_{q}-\mathrm{Id})\,f\equiv 0,$ thereby correcting the lifts to genuine $GL(q)$-invariants. In parallel, within the kernel itself we perform a weightwise computation of $\Sigma_{q}$- and $GL(q)$-invariants and then apply a largest-weight correction to non-zero $GL(q)$-invariants. Below we construct in detail our algorithm as sketched.
$\displaystyle\trianglerightIntegers$ $q\geq 1$, $n\geq 0$ with $n\equiv q\pmod{2}$; base field $\mathbb{F}_{2}$.
@@
$\displaystyle\trianglerightfunction$ $\Sigma_{q}/GL(q)$-On-Kernel-Weight($\mathsf{DS},\;\ker L,\;\mathcal{I}_{w}$)
@@
$\displaystyle\triangleright$  ($\Sigma_{q}$-stage) Build stacked matrix of $(\rho_{j}-\mathrm{Id})\sum_{i}\gamma_{i}u_{i}$ for $j=1,\dots,q-1$; find nullspace
$\displaystyle\triangleright$  Obtain $\Sigma_{q}$-basis $\{\sum_{i}\gamma^{(t)}_{i}u_{i}\}_{t}$
$\displaystyle\triangleright$  ($GL(q)$-stage) Build matrix of $(\rho_{q}-\mathrm{Id})\sum_{t}\beta_{t}(\Sigma\text{-basis})_{t}$; find nullspace
@@
$\displaystyle\triangleright$  return ($\Sigma_{q}$-basis, $GL(q)$-basis) in this weight
@@
$\displaystyle\triangleright$   $(Sigma_{q}[\omega],\,GL(q)[\omega])\leftarrow$ $\Sigma_{q}/GL(q)$-On-Kernel-Weight($\mathsf{DS\_src},\ker L,\mathcal{I}_{\omega}$)
@@
$\displaystyle\triangleright$  [Step 4] Correction inside kernel (largest weight with $GL(q)\neq 0$):
$\displaystyle\triangleright$  if all $GL(q)[\omega]$ are empty then report $GL(q)$-invariants in kernel $=0$
@@
$\displaystyle\triangleright$   pick $\omega^{\star}=\max\{\omega:GL(q)[\omega]\neq\emptyset\}$; set $\mathcal{L}:=\{i\in\mathcal{K}:\text{weight}(i)<\omega^{\star}\}$
$\displaystyle\triangleright$   for each $g_{\max}\in GL(q)[\omega^{\star}]$ do
@@
    Bit-packed matrices and Gaussian elimination over $\mathbb{F}_{2}$. All nullspace computations (Kameko matrix, $\Sigma_{q}$-stage, $GL(q)$-stage) use bit-packed matrices, so elimination and back-substitution become word-level XOR operations. This is cache-friendly and significantly faster than dense arithmetic over $\mathbb{F}_{2}$.